(VI) QUESTIONS:

(i) State the relationship between pressure and volume as stated by Boyle's law. Remember to state the condition of temperature.

(iii) If the pressure of a sample of gas in a cylinder is 760 mm of Hg (1 atom) while occupying a 10-mL (milliliters) volume at 32°C. What is the pressure due to the sample of gas when the cylinder is compressed down to 5 milliliters (mL)? (iii) Consider a balloon with volume 22.4 liter and pressure 1 atom at some constant temperature. What will be new volume if pressure is doubled to 2.0 atom?
Answer #1

I) for T= constant

PV = constant

II) as volume gets halved , pressure becomes double

P = 2× 760 = 1520 mm of Hg

III) as pressure is doubled , volume gets halved

V = (1/2) × 22.4 = 11.2 litre